OpenStudy (anonymous):

(-7x-6)(-x^2+6x-2) - Find the Product.

OpenStudy (mathstudent55):

Do you have an idea of how to start?

OpenStudy (skullpatrol):

You need an idea to get an answer, right?

OpenStudy (skullpatrol):

$$\huge (-7x-6)(-x^2+6x-2) $$

OpenStudy (skullpatrol):

Recall the distributive property: $$\huge a(b+c)=ab+ac$$

OpenStudy (skullpatrol):

$$\ a(b+c+d)=ab+ac+ad$$ Suppose $$a=-7x-6$$$$b=-x^2$$$$c=6x$$$$d=-2$$

OpenStudy (skullpatrol):

Also you will need to use the distributive property in the following form: $$\huge (b+c)a=ba+ca$$

OpenStudy (skullpatrol):

@Jdubose2 Are you still there?